Example: In a class of 3 students, the average score on the final exam is 75% and standard deviation of 5%. Amy got 80% on the final exam. How many standard deviations better than the mean is that? Mandy got 70%. How many standard deviations is Mandys score deviate from the mean? Judy got 75%. How many standard deviation is Judys score deviate from the mean?
Answer #1

We are given,

Mean = μ = 75.SD= sigma = 5

Q1) Now, Amy got 80% which can be expressed as

80=75+5=mu +1sigma

Hence, it is 1 standard deviation better than the mean.

Q2)Mandy got 70% which can be express as

70=75-5=mu -1sigma

Thus, Mandy's score deviates from the mean (here, less than the mean) by 1 standard deviation.

Q3) Judy got 75% which is equal to the mean i.e.

75=mu =mupm 0sigma

Judy''s score does not deviate from the mean.